OpenStudy (anonymous):

what rate percent is 1 minute 12 second to an hour ?

OpenStudy (angela210793):

72/3600=1/50

OpenStudy (angela210793):

so =2%

OpenStudy (anonymous):

1min 12 sec = 72 sec and i hour = 3600 sec for rate 72/3600 = 0.02 rate 0.02 *100=2 %

OpenStudy (anonymous):

ok !! thnxx both neel & angela (smarty) :p

OpenStudy (angela210793):

This was ur q? I didn't notice that u have posted this :P:P:P:P

OpenStudy (anonymous):

actually wat r saying i dnt get u angela !! can u say it once more clearly ?

OpenStudy (angela210793):

Nthng...4get it :)

OpenStudy (anonymous):

say it yaar !! i posted (:p) for saying u smarty

OpenStudy (angela210793):

^_^ NTHNG!!!

OpenStudy (anonymous):

okk !1 chill ^_^

OpenStudy (angela210793):

lol :P :)

OpenStudy (anonymous):

hehe !